# Beyond Calibration: Assessing the Probabilistic Fit of Neural Regressors via Conditional Congruence

### v2

@@ B.3 Addressing Memory Limitations
Appendix C ADDITIONAL EXPERIMENTS
C.1 Using CCE for Machine Learning with Rejection
The rising subfield of “Machine Learning with a Reject Option” centers around uncovering principled techniques for avoiding predictions when they are expected to be inaccurate or probabilistically misaligned [56]. A typical setup defines some “unreliability score” that we can compute on an arbitrary test input. We then select a threshold, and for any input with an unreliability score above that threshold, we withhold from making a prediction.
C.2 Understanding the Behavior of MCMD on Known Distributions
To build intuition about the behavior of the MCMD (and illustrate why we use it as the backbone of our CCE metric), we show how it estimates distance in a variety of settings where the data generating distribution is known. We simulate two data-generating processes, one heteroscedastic, one homoscedastic. For each, we sample 1000 $(x,y)$ pairs to represent the ground truth. We then draw 500 $(x^{\prime},y^{\prime})$ pairs from 1) the same distribution; 2) a distribution with the same mean but smaller variance; 3) a distribution with the same mean but higher variance; 4) a distribution with a different mean and smaller variance; 5) a distribution with a different mean and higher variance; and 6) a distribution with an entirely different conditional relationship. Pairing each of (1-6) with the original distribution, we compute and plot the MCMD ($\sqrt{\widehat{\text{MCMD}^{2}}}$) between $(x,y)$ and $(x^{\prime},y^{\prime})$ across a uniform grid spanning the input space as in [33]. In addition to the MCMD plot, we report the mean value of the MCMD. We compute the MCMD with the regularization parameter $\lambda=0.1$. To measure similarity in both $x$ and $y$, we employ the RBF kernel $k_{\gamma}(\mathbf{x},\mathbf{x^{\prime}})=\exp{(-\gamma||\mathbf{x}-\mathbf{x^{\prime}}||^{2})}$, with $\gamma=1$ for $k_{\mathcal{X}}$, and $k_{\mathcal{Y}}$ selected as described in Section 3.2. Results are visualized in Figures 8 and 9.
Overall, the MCMD appears to behave quite intuitively, with higher values consistently corresponding to areas where the distributions vary from each other more in both mean and standard deviation. Beyond the “eye test”, we specifically highlight a few key desirable characteristics of this metric. First, we see from columns 4 and 5 of Figures 8 and 9 that poor mean fit is penalized more harshly by MCMD when the corresponding variance is low. This indicates that the MCMD requires high variance to explain misalignment of means, which is arguably the correct stipulation to enforce. Additionally, from columns 2 and 3, we observe that the MCMD recognizes the case where the means match but the spread of two distributions is different. This suggests the MCMD is robust at identifying various forms of potential discrepancy between a predictive distribution and the ground truth.
C.3 Hyperparameter Study
(a) The effect of $\lambda$ on MCMD computations over synthetic data.
Figure 10: Studying the impact of the regularizer $\lambda$.
Computing the CCE for a given model requires a handful of up-front decisions — primarily, which functions to use for $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$ (including kernel hyperparameters) and what values to select for $\lambda$ and $\lambda^{\prime}$. Additionally, one must decide how many Monte Carlo samples, $\ell$, to take when forming $\mathcal{S}^{\prime}$ from a model $f$. In Section 3.2, we provide a variety of “default” settings that have proven effective in our work. To further investigate the impact of the choice of kernel, $\lambda$, and $\ell$, we explore the empirical impact of these hyperparameters. To facilitate our study, we compare samples from two data-generating distributions via MCMD: $(X,Y)$, where $X\sim\texttt{Uniform}(0,2\pi),\text{ }Y|X\sim\mathcal{N}(\cos(X),\frac{1}{4})$ and $(X^{\prime},Y^{\prime})$, where $X^{\prime}\sim\texttt{Uniform}(0,2\pi),\text{ }Y^{\prime}|X^{\prime}\sim\mathcal{N}(X^{\prime}-5,\frac{1}{4})$. We also run a series of CCE computations (with differing hyperparameters) for models trained on COCO-People.
C.3.1 Investigating $\lambda$
Figure 11: The effect of the number of Monte Carlo samples, $\ell$, on $\overline{\mbox{CCE}}$ for three models trained on AAF.
To isolate the impact of $\lambda$, we first focus on how it affects the MCMD computation between our samples from $(X,Y)$ and $(X^{\prime},Y^{\prime})$ as specified above. We fix our kernels $k_{\mathcal{X}}$ (RBF kernel with $\gamma=0.5$) and $k_{\mathcal{Y}}$ (RBF kernel with $\gamma=\frac{1}{2\hat{\sigma^{2}_{Y}}}$). We vary $\lambda$ along a grid of values between 0 and 1, then compute the MCMD across the input space. For simplicity, we set $\lambda^{\prime}=\lambda$. In Figure 10(a) we plot these MCMD values (second row), along with a scatterplot of the samples being compared (first row). Note that when $\lambda$ is very small, the MCMD becomes quite sensitive to individual points (observe, for example, the spike in MCMD around $2\pi$ for $\lambda=0.001$). As $\lambda$ increases, the MCMD between the two samples grows smoother and tends toward zero. This behavior agrees with [33]’s categorization of $\lambda$ as a regularizer ($\lambda$ controls the regularization term in the empirical loss [33] minimize to obtain their estimate of MCMD). Thus, we can consider $\lambda$ as a knob that we can turn depending on how much we “believe” that the samples represent the underlying conditional distributions being compared. When $\lambda$ is large, we discard all evidence from our sample that the distributions are different (thus, the MCMD is 0). When $\lambda$ is small, we place more weight on the individual deviations observed between samples. Generally, $\lambda$ values around 0.05 or 0.1 exhibit reasonable behavior. We set $\lambda=0.1$ in our experiments in the main body of the paper.
We also investigate the effect of $\lambda$ on CCE (which calculates MCMD using samples obtained from a model) for three neural networks trained on the COCO-People dataset. Using the same regressors discussed in Section 4.2.1, we compute $\overline{\mbox{CCE}}$ across a grid of $\lambda$ values for each model and display the results in Figure 10(b). We observe a similar pattern as before: increasing $\lambda$ shrinks the mean CCE values towards zero, and we aggressively discard evidence from the data. Interestingly, the relative ordering of the three models is generally preserved across the domain of $\lambda$, suggesting that using CCE to make comparisons across models is fairly robust to the choice of $\lambda$.
C.3.2 Modifying $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$
We next investigate the impact of the $\gamma$ hyperparameter in $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$, which is present in both the RBF and Laplacian kernels. Specifically, we fix $\lambda=0.1$ and vary $\gamma$ simultaneously in $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$. We perform this experiment on the synthetic samples from $(X,Y)$ and $(X^{\prime},Y^{\prime})$ we generated earlier. We use the RBF kernel, but since $\gamma$ induces a similar effect on the Laplacian kernel, we expect an equivalent analysis. Figure 12(a) displays how MCMD computations are affected as $\gamma$ increases. Since the RBF kernel essentially computes similarities by centering a Gaussian at each point with standard deviation $\frac{1}{\sqrt{2\gamma}}$, we translate each $\gamma$ value into an “ellipse of influence” spanning two standard deviations in both the $x$ and $y$ direction. These are visualized in light red around each point in the dataset. As $\gamma$ increases, the number of points that are considered similar to a given $(x,y)$ pair decreases (since the implied Gaussian becomes more concentrated). This increases the sensitivity of the MCMD computation, similarly to selecting low values of $\lambda$. Though most practical applications of MCMD (such as our CCE computation) will not use the RBF kernel for the input values, this behavior provides a good intuition for selecting $\gamma$ for $k_{\mathcal{Y}}$. A workable heuristic is to select $\gamma=\frac{1}{2\hat{\sigma^{2}_{Y}}}$, which ensures that the majority of similarities in $\mathcal{Y}$ are nonzero, with closer points exerting more influence.
(a) The effect of the kernel hyperparameter $\gamma$. Red ellipses indicate the region around each point where similarity (as implied by the RBF kernel) is meaningfully greater than zero.
(b) The effect of the choice of kernel on the MCMD computation.
Figure 12: Studying the impact of the kernel functions $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$ on the MCMD.
Lastly, we directly compare MCMD values obtained with three different kernels: RBF, Laplacian, and Polynomial (applying the same kernel class to the input and the output spaces). For the RBF and Laplacian kernels, we set $\gamma=2$ for $k_{\mathcal{X}}$ and $\gamma=\frac{1}{2\hat{\sigma^{2}_{Y}}}$ for $k_{\mathcal{Y}}$. For the polynomial kernel, we use a degree of 3, offset of 1, and multiply the inner product by $0.02$ (mostly to ensure MCMD values live on the same scale). For all MCMD computations, we set $\lambda=0.1$. Figure 12(b) displays how the MCMD differs across $\mathcal{X}$ for each kernel. We note that each method largely agrees as to the regions with the most and least discrepancy between the samples. Interestingly, the polynomial kernel produces the smoothest MCMD curve, followed by RBF and Laplacian. From these results, we suspect that the choice of kernel does not matter as much as the hyperparameters associated with that kernel.
C.3.3 Increasing the Value of $\ell$
Finally, we examine the impact of the number of Monte Carlo samples, $\ell$, on $\overline{\mbox{CCE}}$ in Figure 11. Specifically, we examine the same three models from Section 4.2.1 and their performance on the AAF dataset. For each model we vary the number of samples, $\ell\in\{1,2,3,4,5\}$. We compute $\overline{\mbox{CCE}}$ three times and plot the mean and standard deviation over the three trials. Importantly, we observe that CCE is robust to the number of Monte Carlo samples; this effect is consistent across models. We recommend setting $\ell=1$ to reduce computational complexity, and do so in all of our experiments.
[…]
Figure 14: Comparison of NLL with CCE for pointwise analysis of probabilistic fit. From left to right: A t-SNE projection of the test split of AAF (color indicates age), the NLL values achieved by the DDPN model, and the CCE values on each test input. NLL is sensitive to outliers and is even undefined for some values (where the probability of the true target was zero). Meanwhile, CCE illuminates the input regions where the model is least congruent. In this case, these regions appear to line up with images of older faces, suggesting a potential inability of the model to reliably assess the ages of elderly people.
@@ Appendix D SHORTCOMINGS OF EXISTING METRICS
D.1 On Bias in the ECE Against Discrete Models
While writing this paper, we noticed some strange behavior with [22]’s ECE score. This led us to investigate further, and we have uncovered significant issues with using this metric to quantify the calibration of discrete regression models. We first restate the definition of ECE in Equation 4:
@@ D.2 On the Limited Applicability of NLL for Point-Wise Analysis
While the ECE is strictly defined in the aggregate sense, it is possible to evaluate the NLL at specific inputs (provided a label is available) to gain some understanding about point-wise probabilistic fit. In practice, however, we find that because NLL lacks context outside the specific point being evaluated, it is sensitive to outliers and produces a rather discontinuous loss surface. Additionally, since NLL has no fixed lower bound, it can be hard to interpret just how close to perfectly-fit a given predictive distribution is ([34] discusses this in some detail). These shortcomings do not extend to the CCE, as its kernelized computation incorporates similar inputs and outputs when assessing congruence and its minimum is 0. As an added bonus, the CCE can be computed on inputs that do not have a label (see Section 4.3 for an explanation of how this works).
To make a more concrete comparison between the NLL and CCE as tools for point-wise comparison, we assess the conditional distributions learned by the DDPN [16] trained on AAF in two ways: by computing the NLL for each test point, and by computing the CCE for each test point (using the validation split as $\mathcal{S}$). We use t-SNE [58] to project the embeddings of each input $\varphi(x_{i})$ to two dimensions for easy visualization, and create a contour plots indicating NLL and CCE across the input space. To build intuition around the geometric structure of the input space, we also provide a plot showing the magnitude of the labels for each test input. Results are displayed in Figure 14.
@@ E.1 AAF
Additional case studies in the style of Section 4.3 for the various models we trained on the AAF dataset can be found in Figures 15, 16, 17, 18, 19, and 20 (excepting the $\beta$-Gaussian, which is displayed in Figure 4 in the main body).
@@ E.2 COCO-People
We also provide additional case studies showing the best and worst examples in terms of CCE for each model trained on the COCO-People dataset (excepting the Poisson that we showcase in Figure 5 in the main body). See Figures 21, 22, 23, 24, 25, and 26.
0.0353
0.0366
0.0371
0.0373
0.0374
0.0953
0.0959
0.0963
0.0964
0.0978

### v3

@@ Appendix D Shortcomings of existing metrics
In this section, we elaborate further on shortcomings of Expected Calibration Error (ECE) and Negative Log Likelihood (NLL).
Figure 14: We simulate four data generating processes for $(X,Y)$ with identical first and second moments. Data are drawn from a Gaussian, Poisson, Negative Binomial, and Double Poisson [12] distribution, where $X\sim\texttt{Uniform}(1,10)$ and $Y|X$ is defined, respectively, as $\texttt{Gaussian}(X,X)$, $\texttt{Poisson}(X)$, $\texttt{NegativeBinomial}(\lceil\frac{X^{2}}{\varepsilon}\rceil,1-{\varepsilon})$, where ($\varepsilon$ << 1), and $\texttt{Double Poisson}(X,1)$. We draw 2000 samples from each distribution. (Row 1) A perfectly specified probabilistic model with mean and 95% centered credible intervals. (Row 2) A histogram of each model’s Probability Integral Transform (PIT). (Row 3) A reliability diagram in the style of Kuleshov et al. [26], along with each model’s ECE ($\alpha=1$). (Row 4) The Conditional Congruence Error (CCE) between samples drawn from each predictive distribution and the ground truth dataset. Although the histograms, reliability diagrams, and ECE values identify the well-specified Gaussian model, they fail to capture calibration for the discrete models. Meanwhile, our proposed approach using CCE correctly indicates that each model is (with some sampling error) perfectly congruent, and thus calibrated.
D.1 On the bias in ECE against discrete models
While writing this paper, we noticed some strange behavior with Kuleshov et al. [26]’s ECE score. This led us to investigate further, and we have uncovered significant issues with using this metric to quantify the calibration of discrete regression models. We first restate the definition of ECE in Equation 4:
@@ D.2 On the limited applicability of NLL for point-wise analysis
While the ECE is strictly defined in the aggregate sense, it is possible to evaluate the NLL at specific inputs (provided a label is available) to gain some understanding about point-wise probabilistic fit. In practice, however, we find that because NLL lacks context outside the specific point being evaluated, it is sensitive to outliers and produces a rather jagged loss surface. Additionally, since NLL has no fixed lower bound, it can be hard to interpret just how close to perfectly-fit a given predictive distribution is (Błasiok et al. [5] discusses this in some detail). These shortcomings do not extend to the CCE, as its kernelized computation incorporates similar inputs and outputs when assessing congruence and its minimum is 0. As an added bonus, the CCE can be computed on inputs that do not have a label (see Section 4.7 for an explanation of how this works).
In the main body of the paper, to make a more concrete comparison between the NLL and CCE as tools for point-wise analysis, we assess the conditional distributions learned by a DDPN [46] trained on AAF in two ways: by computing the NLL for each test point, and by computing the CCE for each test point (using the validation split as $\mathcal{S}$). We use t-SNE [43] to project the embeddings of each input $\varphi(x_{i})$ to two dimensions for easy visualization, and create a contour plots indicating NLL and CCE across the input space. To build intuition around the geometric structure of the input space, we also provide a plot showing the magnitude of the labels for each test input. Results are displayed in Figure 4, and related discussion can be found in Section 4.6.
@@ Appendix E Additional case studies
In this section, we provide additional case studies (in the style of Section 4.7) showcasing how CCE can be used to diagnose specific input regions in which a model’s probabilistic outputs are unreliable.
@@ E.1 AAF
Case studies for the various models we trained on the AAF dataset can be found in Figures 15, 16, 17, 18, 19, and 20 (excepting the $\beta$-Gaussian, which is displayed in Figure 5 in the main body).
@@ E.2 COCO-People
We also provide case studies showing the best and worst examples in terms of CCE for each model trained on the COCO-People dataset. See Figures 21, 22, 23, 24, 25, and 26.
0.0255 (0.009)
0.0266 (0.012)
0.0270 (0.010)
0.0271 (0.009)
0.0295 (0.013)
0.1769 (0.027)
0.1761 (0.028)
0.1701 (0.031)
0.1675 (0.032)
0.1634 (0.023)